OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 9x^2 - 18x - 45 = 0

OpenStudy (anonymous):

divide by 9

OpenStudy (anonymous):

done.

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

(x-9)^2=11 but i think thats wrong I'm starting over again.

OpenStudy (anonymous):

can we do it step by step please?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

and yes that is wrong lets start with \[9x^2-18x-45=0\] and divide all by \(9\) to make the leading coefficient one just divide each number by \(9\) and tell me what you get

OpenStudy (anonymous):

i mean not as a final answer, just what you get when you divide all by 9

OpenStudy (anonymous):

9x^2-2x-5=0

OpenStudy (anonymous):

you forgot the most important part you are trying to make the leading coefficient (the number in front of \(x^2\) in to \(1\)

OpenStudy (anonymous):

let me know when you get \[x^2-2x-5=0\]

OpenStudy (anonymous):

lol wow i feel dumb lol. i thought i divided wrong. i got it now.

OpenStudy (anonymous):

then add \(5\) to both sides

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

x^2-2x=5

OpenStudy (anonymous):

ok now what is half of \(2\) ?

OpenStudy (anonymous):

(x-1)^2=6

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

yep. thankyou:)